I have a problem with solving this question..

20cm3 of methane was released with 18cm3 of oxygen and it exploded.Calculate the volume of the resulting gas after the reaction mixture was cooled to room temperature.
1
Expert's answer
2020-06-15T13:56:29-0400

The explosion and burning of methane with oxygen results in the formation of carbon dioxide and water according to the following reaction:

CH4 + 2O2 = CO2 + 2H2O

Next, the number of moles of methane and oxygen must by calculated. As 1 mol of gas occupies 22.4 L at standard conditions, the numbers of moles of methane and oxygen are:

n(CH4) = V(CH4) / 22.4 L = 20 cm3 / 22.4 L = 0.02 L / 22.4 L = 8.9 × 10-4 mol

n(O2) = V(O2) / 22.4 L = 18 cm3 / 22.4 L = 0.018 L / 22.4 L = 8.04 × 10-4 mol

As n(O2) < n(CH4), oxygen is a limiting reagent. As a result, it must be used to calculate the number of moles of carbon dioxide produced during the reaction. As 1 mole of CO2 can be produced from two moles of oxygen:

n(CO2) = n(O2) / 2 = 8.04 × 10-4 mol / 2 = 4.02 × 10-4 mol

Finally, as 1 mol of gas occupies 22.4 L, carbon dioxide occupies:

V(CO2) = n(CO2) × 22.4 L = 4.02 × 10-4 mol × 22.4 L = 9 × 10-3 L = 9 cm3


Answer: 9 × 10-3 L or 9 cm3